OpenStudy (anonymous):

how would you find the quotient and remainder when 3x^4-x^2 is divided by x^3-x^2+2. I know we do long division but can someone help me go thought steps PLEASE!

OpenStudy (whpalmer4):

Have you tried the problem yet?

OpenStudy (whpalmer4):

Look at the highest order term in both the dividend and the divisor. They are respectively \(3x^4\) and \(x^3\). What is the largest factor you can multiply the latter by without exceeding the former?

OpenStudy (anonymous):

well the quotient be 3x+3 and remainder be 2x^2-6x-6??

OpenStudy (whpalmer4):

Well, we can check that: the quotient * the divisor + the remainder should give us the dividend. \[(3x+3)(x^3-x^2+2) = 3x^4 -3x^3 + 6x + 3x^3 -3x^2+6 = 3x^4-3x^2+6x+6\] \[3x^4-3x^2+6x+6 - (2x^2-6x-6) = 3x^4-3x^2+6x+6-2x^2+6x+6=\] \[3x^4-5x^2+12x+12\] Does that look like the right answer?

OpenStudy (anonymous):

no i guess it doesn't look right

OpenStudy (whpalmer4):

Hmm, but it is right, where did I make a mistake?

OpenStudy (whpalmer4):

\[(3x+3)(x^3-x^2+2) = 3x^4-3x^2+6x+6\] Oh, I subtracted the remainder, instead of adding it. Oops. \[3x^4-3x^2+6x+6 + 2x^2-6x-6 = 3x^4 -3x^2 + 2x^2 + 6x - 6x + 6 - 6 =\]\[ 3x^4-x^2\]

OpenStudy (anonymous):

oh okay thank you super much!

OpenStudy (whpalmer4):

Sorry about the wrong turn there! I was still thinking of doing the division, I guess.

OpenStudy (anonymous):

Its okay! thank you for helping!!

OpenStudy (whpalmer4):

Got another one?

OpenStudy (anonymous):

yes but should i make a new question or should i post here?

OpenStudy (whpalmer4):

New question is always good, I think.